OpenStudy (anonymous):

Use the quadartic formula to solve the equation below. 2x^2+13x=4

OpenStudy (anonymous):

Do you mean quadratic? Use the formula and plug in the coeff of the x^2 for a, the coeff of the x term for b and the coeff of the constant term for c. 2x^2+13x=4 ---> 2x^2+13x-4 2x^2--->a=2 13x------>b=13 -4---->c=-4

OpenStudy (anonymous):

-b + or - sqrt(b^2-4ac) all divided by 2a

OpenStudy (anonymous):

Hold on i have chocies

OpenStudy (anonymous):

it would be better to say plus and minus, you get two answers

OpenStudy (anonymous):

they should both work as solutions to the equation

OpenStudy (anonymous):

If you're doing homework they probably want both answers.

OpenStudy (anonymous):

A.) x= 13+√201 -------------- 2 B.)x=-13+√177 ---------------- 4 C.) x= -13+√ 201 ---------------- 2 D.) x= -13 +√ 201 ----------- 4

OpenStudy (anonymous):

AND + HAS a line under it \[\pm\]

OpenStudy (anonymous):

Close, not quite.

OpenStudy (anonymous):

substitute your values in the quadratic formula and see which is the correct one.

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

what's the difference between C and D

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

c has a 2 and d has a 4

OpenStudy (anonymous):

Well the answer is exactly the same as D, only it's + and - sqrt201

OpenStudy (anonymous):

katie, there is no point guessing the answer. take some time and substitute your vlues into the quadratic equation. that will give you the correct answer.

OpenStudy (anonymous):

so its D?

OpenStudy (anonymous):

If the + is a + - line thing, yes

OpenStudy (anonymous):

±

OpenStudy (anonymous):

yes it has this \[\pm\] under it so its D

OpenStudy (anonymous):

Yeah, i mean it's exactly the same as the formula, it just matters that you plug in the right numbers.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Check your work on http://www.wolframalpha.com/input/?i=2x^2%2B13x%3D4+quadratic+formula

OpenStudy (anonymous):

I've gotta go, goodluck

OpenStudy (anonymous):

Thanks